\begin{thebibliography}{}

\bibitem{BalD} 
J.\ S.\ Balakrishnan and N.\ Dogra
\emph{Quadratic Chabauty and rational points I: p-adic heights},
\textbf{arXiv:1601.00388}.


\bibitem{BBM}
A.\ Besser, J.\ Balakrishnan and J.\ S.\ M\"{u}ller,
\emph{Quadratic Chabauty: p-adic height pairings and integral points on hyperelliptic curves}, 
Journal f\"{u}r die reine und angewandte Mathematik \emph{720} (2016), 51--79.


%\bibitem{Bruin} N.\ Bruin,
%\emph{Chabauty methods using elliptic curves},
%J.\ reine angew.\ Math.\  \textbf{562}  (2003), 27--49.

%\bibitem{ChenSiksek}
%I.\ Chen and S.\ Siksek,
%\emph{Perfect powers expressible as sums of two cubes},
%Journal of Algebra \textbf{322} (2009), 638--656.


%\bibitem{Chabauty} C.\ Chabauty,
%{\em Sur les points rationnels des vari\'{e}t\'{e}s
%alg\'{e}briques dont l'irr\'{e}gularit\'{e}
%est sup\'{e}rieure \`{a} la dimension},
%C.\ R.\ Acad.\ Sci.\ Paris {\bf 212} (1941), 1022--1024.

\bibitem{Coleman} R.\ F.\ Coleman,
{\em Effective Chabauty},
Duke Mathematical Journal {\bf 52} (1985), No.\ 3, 765--770.


%\bibitem{modcurves} J.\ Cremona, J.-C.\ Lario, J.\ Quer, K.\ Ribet (editors), {\em Modular Curves and Abelian Varieties}, Progr.\ Math.\ {\bf 224}, Birkh\"{a}user, Basel, 2004.  



\bibitem{Fa1} G.\ Faltings,
\emph{Endlichkeitss\"{a}tze f\"{u}r abelsche Variet\"{a}ten \"{u}ber 
Zahlk\"{o}rpern}, Invent.\ Math.\ \textbf{73} (1983), no.\ 3, 349--366.

%\bibitem{Fa2} G.\ Faltings,
%\emph{Diophantine approximation on abelian varieties}, Ann.\ of Math. (2) \textbf{133} (1991),
%no. 3, 549--576.

%\bibitem{FW} E.\ V.\ Flynn and J.\ L.\ Wetherell,
%\emph{Finding rational points on bielliptic genus 2 curves},
%Manuscripta Math.\ \textbf{100} (1999), no.\ 4, 519--533.


%\bibitem{FHS} N.\ Freitas, B.\ Le Hung and S.\ Siksek,
%\emph{Elliptic curves over real quadratic fields are modular},
%Inventiones Mathematicae \textbf{201} (2015), 159--206.

%\bibitem{Frey} G.\ Frey,
%\emph{Curves with infinitely many points of fixed degree},
%Israel Journal of Mathematics \textbf{85} (1994), 79--83.

%\bibitem{FS} N.\ Freitas and S.\ Siksek,
%\emph{The asymptotic Fermat's Last Theorem for five-sixths of real quadratic fields},
%Compositio Mathematica, published online 6/3/2015.

%\bibitem{FermatII} N.\ Freitas and S.\ Siksek,
%\emph{Fermat's Last Theorem for some small real quadratic fields},
%Algebra \& Number Theory 
%\textbf{9} (2015), 875--895.

%\bibitem{irred} N.\ Freitas and S.\ Siksek,
%\emph{Criteria for irreducibility of mod $p$ representations of Frey curves}, 
%to appear in the Journal de Th\'{e}orie des Nombres de Bordeaux.

\bibitem{Giovanetti}
F.\ Giovanetti,
\emph{On the N\'eron--Severi group of Jacobians of curves with automorphisms},
Indag.\ Mathem.\ \textbf{4} (1993), 311--320.


%\bibitem{GR}
%B.\ H.\ Gross and D.\ E.\  Rohrlich,
%{\em Some results on the Mordell-Weil group of the Jacobian of the Fermat curve},
%Invent.\ Math.\ {\bf 44} (1978), 201--224.


%\bibitem{Kalyanswamy}
%S.\ Kalyanswamy,
%\emph{Remarks on Automorphy of Residually Dihedral Representations},
%\texttt{arXiv:1607.04750}.

%\bibitem{Kamienny}
%S.\ Kamienny,
%\emph{Torsion points on elliptic curves and $q$-coefficients of modular forms},
%Invent.\ Math.\ \textbf{109} (1992), 221--229.

%\bibitem{KRZ}
%E.\ Katz, J.\ Rabinoff and
%D.\ Zureick-Brown, 
%\emph{Uniform bounds for the number of rational points 
%on curves of small Mordell--Weil rank}, 
%Duke Mathematical Journal, to appear.


%\bibitem{KW1}
%C.\ Khare and J.-P.\ Wintenberger,
%\emph{Serre's modularity conjecture. I},
% Invent.\ Math.\ \textbf{178} (2009), no.\ 3, 485--504.

%\bibitem{KW2}
%C.\ Khare and J.-P.\ Wintenberger,
%\emph{Serre's modularity conjecture. II},
% Invent.\ Math.\ \textbf{178} (2009), %no. 3,
%505--586.

\bibitem{Kim}
M.\ Kim,
\emph{The unipotent Albanese map and Selmer varieties for curves}, 
Publ.\ Res.\ Inst.\ Math.\ Sci.\ \textbf{45} (2009), no.\ 1, 89--133.


%\bibitem{KolyvaginLogachev}
%V.\ A.\ Kolyvagin and D.\ Yu.\ Logachev, 
%\emph{Finiteness of the Shafarevich--Tate group
%and the group of rational points on some modular abelian varieties}, 
%Leningrad Math.\ J.\ \textbf{1}(1990), 1229--1253.


%\bibitem{Levi} B.\ Levi, 
%\emph{Sull’equazione indeterminata del $3^{o}$ ordine}, Atti del IV Congresso Internazionale
%dei matematici, Roma 1908 \textbf{2}, 175--177.

\bibitem{MP} W.\ McCallum and B.\ Poonen,
\emph{The method of Chabauty and Coleman},
pages 99--117 of \emph{Explicit Methods in Number Theory},
ed.\ K.\ Belabas et al., Panoramas et Synth\`{e}ses \textbf{36},
SMF, 2012.

%\bibitem{Maz77} B.\ Mazur,
%\emph{Modular curves and the Eisenstein ideal},
%Publ. Math. IHES (1977) \textbf{47}, 33--186.


%\bibitem{Mazur2} B.\ Mazur,
%\emph{Rational Points on Modular Curves},
%pages 107--147 of 
%\emph{Modular Functions of One Variable V}, 
%J.-P.\ Serre and D.\ B.\ Zagier (editors), 
%Springer-Verlag, 1977.


%\bibitem{Merel} L.\ Merel,
%\emph{Bornes pour la torsion des courbes elliptiques sur les corps de nombres},
%Invent.\ Math.\ {\bf 124} (1996), 437--449.

%\bibitem{MSS}
%J.\ Merriman, S.\ Siksek and N.\ Smart,
%\emph{Explicit $4$-descents on an elliptic curve}, 
%Acta Arithmetica \textbf{LXXVII.4} (1996), 358--404.  

%\bibitem{Mourao} M.\ Mourao,
%\emph{Extending elliptic chabauty to higher genus curves},
%Manuscripta Mathematica \textbf{143} (2014), 355--377.

%\bibitem{Park} J.\ Park,
%\emph{Effective Chabauty for symmetric powers of curves},
%\texttt{arXiv:1606.05195}.

\bibitem{Pyle} E.\ E.\ Pyle,
\emph{Abelian varieties over $\Q$ with large endomorphism algebras
and their simple components over $\overline{\Q}$},
Progress in Mathematics \textbf{24} (2004), 189--239.

%\bibitem{Ribet}
% K.\ A.\ Ribet,
%{\em On modular representations of $\Gal(\overline{\Q}/\Q)$
%arising from modular forms},
%Invent.\ Math.\ {\bf 100} (1990), 431--476.

\bibitem{Rohrlich} D.\ E.\ Rohrlich,
\emph{Modular curves, Hecke correspondences, and $\mathrm{L}$-functions},
pages 41--100 of 
\emph{Modular Forms and Fermat's Last Theorem}
(G.\ Cornell, J.\ H.\ Silverman
and G.\ Stevens eds),
Springer-Verlag, 1997.



\bibitem{Sikiii} S.\ Siksek,
\emph{Chabauty and the Mordell--Weil Sieve}, 
pages 194--224 of \emph{Advances on Superelliptic Curves and Their 
Applications}, (L.\ Beshaj, T.\ Shaska, E.\ Zhupa eds), IOS Press, 2015. 

\bibitem{Stein} W.\ A.\ Stein,
{\em Modular Forms: A Computational Approach},
American Mathematical Society, 2007.



\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb} 
\usepackage{amsmath} 
\usepackage{amscd}
\usepackage{amsbsy}
\usepackage{comment}
\usepackage[matrix,arrow]{xy}

\DeclareMathOperator{\End}{End}
\DeclareMathOperator{\rk}{rank}

\DeclareMathOperator{\NS}{NS}

\DeclareMathOperator{\GL}{GL}
\DeclareMathOperator{\SL}{SL}
\DeclareMathOperator{\Sel}{Sel}

\DeclareMathOperator{\rank}{rank}
\DeclareMathOperator{\Pic}{Pic}

\DeclareMathOperator{\Gal}{Gal}

\newcommand{\Q}{{\mathbb Q}}
\newcommand{\Z}{{\mathbb Z}}
\newcommand{\cA}{\mathcal{A}}

\newcommand{\OO}{{\mathcal O}}

\begin {document}

\newtheorem{thm}{Theorem}
\newtheorem{lem}{Lemma}[section]

\newtheorem{cor}[lem]{Corollary}

\theoremstyle{definition}

\theoremstyle{remark}

\title[]{Quadratic Chabauty for Modular Curves}
\author{Samir Siksek}
\address{Mathematics Institute\\
	University of Warwick\\
	Coventry\\
	CV4 7AL \\
	United Kingdom}

\email{s.siksek@warwick.ac.uk}

\date{\today}
\thanks{
The author is supported by
the EPSRC {\em LMF: L-Functions and Modular Forms} Programme Grant
EP/K034383/1.
}

\subjclass[2000]{Primary 11G30}

\begin {abstract}
Let $X/\Q$ be a curve of genus $g \ge 2$ with Jacobian $J$ and let $\ell$ be a prime
of good reduction. Using Selmer varieties, Kim defines a decreasing sequence
\[
X(\Q_\ell) \supseteq X(\Q_\ell)_1 \supseteq X(\Q_\ell)_2 \supseteq \cdots 
\]
all containing $X(\Q)$.
Thanks to the work of Coleman, the \lq Chabauty set\rq\ $X(\Q_\ell)_1$ is known
to be finite provided the \lq Chabauty condition\rq\ $\rank J(\Q) < g$ holds.
In this case one has a practical strategy that often succeeds in computing the set of rational points $X(\Q)$.
Balakrishnan and Dogra have recently shown that the \lq quadratic Chabauty set\rq\ $X(\Q_\ell)_2$
is finite provided 
\[
\rank J(\Q) < g + \rank \NS(J)-1,
\]
where $\NS(J)$ is the 
N\'eron-Severi group of $J/\Q$. In view of this it is interesting to give families of 
curves where $\rank \NS(J) \ge 2$ and where therefore quadratic Chabauty is
more likely to succeed than classical Chabauty. In this note we show that this is indeed the 
case for all modular curves of genus $\ge 3$.
\end {abstract}
\maketitle

\section{Introduction}

Let $X$ be a smooth algebraic curve of genus $g \ge 2$ defined over $\Q$ and write
$J$ for its Jacobian. 
Let $\ell$ be a 
prime of good reduction for $X$. The method of Chabauty and Coleman \cite{Coleman} yields
a subset $X(\Q_\ell)_1 \subseteq X(\Q_\ell)$ (which we may term the \lq Chabauty set\rq)  containing $X(\Q)$
that can be explicitly described in terms of Coleman integrals.
Coleman \cite{Coleman} showed that $X(\Q_\ell)_1$ is finite provided
the \lq Chabauty condition\rq\ holds:
\begin{equation}\label{eqn:chab1}
\rank J(\Q) < g .
\end{equation}
This implies that $X(\Q)$ is finite. Of course
Faltings' Theorem \cite{Fa1} asserts the
finiteness of $X(\Q)$ without assumptions beyond $g\ge 2$. 
However, Faltings' Theorem is ineffective. In contrast (see for example
\cite{MP} or \cite{Sikiii})  
the method of Chabauty and Coleman often allows for the computation of $X(\Q)$ 
provided the Chabauty condition \eqref{eqn:chab1} is satisfied
and one knows a Mordell--Weil basis for $J(\Q)$ (or even for a subgroup of
$J(\Q)$ of finite index).

In \cite{Kim}, Kim defined a family of Selmer varieties $\Sel(U_n)$ giving a decreasing sequence of subsets
\[
X(\Q_\ell) \supseteq X(\Q_\ell)_1 \supseteq X(\Q_\ell)_2 \supseteq \cdots \supseteq X(\Q),
\]
which can be computed in terms of iterated Coleman integrals. This offers hope 
of a strategy to compute $X(\Q)$ even if the Chabauty condition \eqref{eqn:chab1} fails.
Indeed, the conjectures of Bloch
and Kato imply that $X(\Q_\ell)_n$ is finite for $n$ sufficiently large \cite{Kim}.
The following theorem of Balakrishnan and Dogra \cite[Lemma 3]{BalD}
gives a  criterion for the finiteness of $X(\Q_\ell)_2$. 
\begin{thm}[Balakrishnan \& Dogra \cite{BalD}]
Suppose
\begin{equation}\label{eqn:chab2}
\rk J(\Q) < g+ \rk \NS(J)-1,
\end{equation}
where $\NS(J)$ is the N\'eron--Severi rank of $J$ over $\Q$. 
 Then $X(\Q_\ell)_2$ is finite.
\end{thm}
Indeed, Balakrishnan \& Dogra 
explain how, under condition \eqref{eqn:chab2}, a function 
can be constructed that 
cuts out $X(\Q_\ell)_2$
explicitly (following \cite{BBM} this method is termed \emph{Quadratic Chabauty}). They then apply this method to determine the rational
points on several genus $2$ curves
that have the form
\begin{equation}\label{eqn:Dogra}
X \; : \; y^2=a_6 x^6+ a_4 x^4+ a_2 x^2+a_0,
\end{equation}
where $\rk J(\Q)=2$. 

The method of Balakrishnan \& Dogra at first seems rather special.
For a \lq generic\rq\ curve $X$ it is known that the N\'eron--Severi
group has rank $1$ (and is in fact spanned by the class of the 
principal polarization). Hence the inequality
\eqref{eqn:chab2} is equivalent to the inequality \eqref{eqn:chab1},
and the method of Balakrishnan \& Dogra does not yield any more than 
the classical
Chabauty. Arguably modular curves are the most interesting family of 
curves. In this note, we give an explicit sufficiency criterion, based on 
the theorem of Balakrishnan \& Dogra, for the finiteness of $X(\Q_\ell)_2$
when $X$ is a modular curve. In fact we show that for any modular curve
$X/\Q$ of genus $\ge 3$ we have $\rank \NS(J) \ge 2$, thus
quadratic Chabauty is more likely to succeed than classical Chabauty.

\section{The Balakrishnan--Dogra Criterion and Modular Curves}

Let $G$ be a congruence subgroup of $\SL_2(\Z)$ and let $X=X_G$ be
the corresponding modular curve which we shall suppose is defined
over $\Q$ (sufficient conditions on $G$ for this to hold are given 
in \cite[Section 1.2]{Rohrlich}). Assume that $X$ has genus $g \ge 2$ and
write $J$ for the Jacobian of $X$.
Let $f_1,\dotsc,f_n$ be representatives
of the Galois orbits of weight $2$ cuspidal eigenforms for $G$ (these may
be computed via the modular symbols algorithm as in \cite{Stein}). 
The theory of Eichler--Shimura attaches to each
$f_i$ an abelian variety of $\GL_2$-type, which we shall denote
by $\cA_i$. 
Then 
\begin{equation}\label{eqn:factor}
J \approx \cA_{1} \times \cdots \times \cA_{n},
\end{equation}
where $\approx$ denotes isogeny over $\Q$.
As $\cA_i$ is of $\GL_2$-type, we have $\End(\cA_i) \otimes \Q$
 is a number field
$F_i$ of degree $\dim \cA_i$, which is either totally real
or CM. The fields $F_i$ are generated by the Hecke eigenvalues of $f_i$.
We arrange the $f_i$ so that $F_1,\dotsc,F_m$ are totally
real and $F_{m+1},\dots,F_n$ are CM.  
\begin{cor}
With the above notation and assumptions,
suppose 
\begin{equation}\label{eqn:betterchab1}
\rank J(\Q)< -1+ 2 \sum_{i=1}^m [F_i:\Q]+\frac{3}{2}\sum_{i=m+1}^n [F_i:\Q].
\end{equation}
Then, for any prime $\ell$ of good reduction for $X$, 
the set $X(\Q_\ell)_2$ is finite.
\end{cor}
\begin{proof}
Note that
\[
g=\dim(J)=\sum_{i=1}^n \dim \cA_i=\sum_{i=1}^n [F_i:\Q].
\]
Thus \eqref{eqn:betterchab1} maybe rewritten as
\begin{equation}\label{eqn:betterchab}
\rank J(\Q)< (g-1)+\sum_{i=1}^m [F_i:\Q]+\frac{1}{2}\sum_{i=m+1}^n [F_i:\Q].
\end{equation}
It is sufficient to show, thanks to the theorem of Balakrishnan \& Dogra,
that
\[
\rank \NS(J) \ge \sum_{i=1}^m [F_i:\Q]+\frac{1}{2}\sum_{i=m+1}^n [F_i:\Q].
\]
We now invoke the isomorphism of $\Q$-vector spaces
\[
\NS(J) \otimes \Q \cong \End^{(s)}(J) \otimes \Q,
\]
where $\End^{(s)}(J)$ is the subring of $\End(J)$ left invariant by
the Rosati involution (see \cite[Section 1]{Giovanetti} for example). 
However, the Rosati involution restricted to $\End(\cA_i)$ induces
complex conjugation on $F_i$ (\cite[page 196]{Pyle}). 
Thus we see that $\End^{(s)}(J) \otimes \Q$
contains the algebra 
\begin{equation}\label{eqn:algebra}
F_1\times \cdots \times F_m \times E_{m+1} \times \cdots \times E_n,
\end{equation}
where, for $m+1 \le i \le n$, we denote the maximal totally real
subfield of $F_i$ by $E_i$. As $F_i$ is CM, we have $[E_i:\Q]=[F_i:\Q]/2$.
The corollary follows.
\end{proof}
It is evident that \eqref{eqn:betterchab} (and hence \eqref{eqn:betterchab1})
is weaker than
Chabauty's condition \eqref{eqn:chab1} whenever $g \ge 3$. If $g=2$
then \eqref{eqn:betterchab} 
is weaker if either $J$ has real multiplication, or is isogenous to 
a product of two elliptic curves.

\medskip

\noindent \textbf{Remark}. If $J$ has multiplicity $1$ (meaning
that the
simple factors $\cA_i$ are pairwise non-isogenous) then 
$\End(J) \cong \End(\cA_1) \times \cdots \times \End(\cA_n)$
and the 
above argument in fact shows that
$\NS(J) \otimes \Q$ is isomorphic to the algebra \eqref{eqn:algebra}.

\end{document}